# The Sparse Frontier: Sparse Attention Trade-offs in Transformer LLMs

### v1

@@ 5 Conclusions
Our study undertakes the most comprehensive comparison of sparse attention methods to date, applied to a range of model scales (up to 72B), sequence lengths (up to 128K), and sparsity levels (up to 95%) across 9 diverse long-sequence tasks. We revealed that, under an isoFLOPS analysis, larger models with high sparsity are more effective than their smaller, dense counterparts, particularly for very long sequences. This finding suggests a shift in strategy where scaling up model size combined with sparse attention mechanisms is crucial for efficiently handling extended sequences.
Another key finding is that the degree of sparsity applicable while statistically ensuring full accuracy retention is on average very high (10-15$\times$) and even increases for larger models during decoding. However, this aggregate perspective obscures a crucial caveat: our analysis reveals that even moderate sparsity levels (e.g., 5x compression) frequently result in significant performance degradation on at least one task for most configurations. This highlights a critical limitation: average performance gains may obscure significant degradation on certain tasks, even at moderate sparsity levels. This task-dependent sensitivity underscores the need for thorough evaluation across diverse benchmarks, covering the full spectrum of potential deployment scenarios. Consequently, sparse attention is not a panacea and always necessitates careful trade-off evaluation, particularly for performance-sensitive applications. Future research should prioritise dynamic sparsity mechanisms that adapt to input and task demands, ideally incorporating performance guarantees.
Zooming in on the results, we observe a trend towards increased performance with greater flexibility and finer granularity in selecting attention interactions for both prefilling and decoding. During prefilling, the optimal sparsification structure (e.g., blocks or verticals and slashes) varies by task, with uniform allocation across layers performing comparably to dynamic allocation. During decoding, page-level Quest excels by preserving the KV cache structure, avoiding the performance degradation associated with token pruning during generation. While more flexible and fine-grained methods incur higher indexing costs, their ability to adapt to diverse task requirements (as evidenced by the differing optimal strategies for prefilling vs. decoding and across tasks) suggests that pursuing such adaptive sparsity mechanisms is a promising direction for future research.
Finally, our study establishes robust scaling laws for sparse attention, which hold true on held-out data, indicating that the observed trends are likely to generalise beyond the tested configurations. Together, these insights demonstrate that sparse attention is bound to play a key role in next-generation LLM architectures, ultimately contributing to more scalable, efficient, and adaptable AI models.
Acknowledgements
This work was supported in part by the UKRI Centre for Doctoral Training in Natural Language Processing, funded by the UKRI (grant EP/S022481/1) and the University of Edinburgh, School of Informatics and School of Philosophy, Psychology & Language Sciences.
@@ A.1 Implementation Details
This section provides supplementary details on the sparse attention patterns evaluated, focusing on hyperparameter tuning and specific configurations used to achieve target compression ratios. We tuned hyperparameters for each pattern using ablation studies on the Qwen-7B model with a 16K sequence length across all tasks, varying compression ratios from 1x to 10x. Our main experiments evaluated performance at fixed compression ratios (1.5x, 2.0x, 2.5x, 3.33x, 5x, 7.5x, 10x, 15x, 20x), using linear interpolation for intermediate values where necessary. Table 6 summarizes the final parameters we used for each pattern, sequence length, and compression ratio.
@@ A.1.1 Block-Sparse Attention
We implement block-sparse attention by dividing the attention matrix into fixed-size blocks. Based on our ablation studies (Figure 6), we selected a block size of 16x16, as smaller blocks consistently yielded better performance. To achieve a target compression ratio, we select the top-$k$ key blocks for each query block, where $k$ is determined via binary search. We always preserve attention sinks (the first key block) and local context (diagonal key blocks corresponding to the query block).
@@ A.1.2 Vertical-Slash Pattern
We implement the Vertical-Slash pattern (Jiang et al., 2024) by allocating a uniform budget to global (vertical columns) and local (slash diagonals) attention components. We select the most important verticals and slashes by approximating attention scores using a limited window of recent query tokens. Our ablation studies (Figure 11) revealed task-dependent optimal approximation window sizes: 512 tokens for retrieval-heavy tasks (Ruler NIAH, Story Retrieval) and 256 tokens for other tasks. This observation correlates with the typical query lengths for these tasks (see Table 5). We consistently preserve the first 4 (prefix) and the most recent 64 (local) tokens. To achieve target compression ratios, we compute the required number of verticals and slashes based on collected attention statistics for each sequence length.
@@ A.1.3 FlexPrefill
We implement FlexPrefill (Lai et al., 2025), which enhances Vertical-Slash by introducing dynamic budget allocation per layer and head, controlled by a coverage parameter $\alpha$ and a minimum budget (min_budget). We set $\tau=0$ in our experiments, hence disabling Query-Aware attention. This choice stemmed from two key considerations: first, our preliminary tests indicated no significant performance gains from enabling it, aligning with the findings reported in the original work; second, this setting isolates the dynamic budget allocation mechanism, allowing us to specifically evaluate its impact compared to the fixed allocation used in the Vertical-Slash pattern. We employ the same task-dependent approximation windows (256 / 512 tokens) and critical token preservation strategy (first 4 prefix, most recent 64 local) as in our Vertical-Slash implementation. Our ablations (Figure 8) indicated that setting min_budget to 512 significantly improved performance, suggesting the importance of maintaining a minimum level of connectivity during prefilling. We achieved target compression ratios by selecting the appropriate $\alpha$ based on attention statistics while keeping min_budget fixed at 512. For high compression ratios where dynamic allocation proved less effective, we set $\alpha=0$, effectively reverting to a uniform allocation of Vertical-Slash.
@@ A.1.4 SnapKV
We implement SnapKV (Li et al., 2024b) by compressing the Key-Value (KV) cache after the prefilling stage and applying a uniform token budget across all heads for the subsequent decoding phase. We predict token importance for decoding by computing attention scores using a window of recent query tokens (approximation window). Our ablations showed an optimal approximation window size of 256 tokens (Figure 10), with no significant task dependency observed, unlike Vertical-Slash and FlexPrefill. We smooth the calculated token importance scores using 1D average pooling with a kernel size of 21 (chosen based on Figure 10). We always preserve the first 4 and the most recent 128 tokens. We control sparsity by setting the ‘token_capacity‘ (token limit per head) to match the target compression ratio.
@@ A.1.5 Ada-SnapKV
We implement Ada-SnapKV (Feng et al., 2024), which extends SnapKV by incorporating dynamic token budget allocation per head. One difference in our implementation between Ada-SnapKV and SnapKV is that we use max-aggregation (instead of averaging) across query positions and heads for score calculation; this empirically proved more effective for adaptive allocation but had no effect for uniform (SnapKV) allocation. We utilize the same smoothing kernel size (21) and critical token preservation strategy (first 4 prefix, most recent 128 local) as in our SnapKV implementation. Our ablations (Figure 8) indicated that providing each head with a minimum budget of 20% of its capacity was optimal. Performance was found to be less sensitive to minimum budget (performing well within the 10-50% range) compared to FlexPrefill’s sensitivity, but degraded sharply when approaching 100% (uniform allocation), underscoring the benefits of dynamic allocation during decoding. We control sparsity by setting ‘token_capacity‘, identically to SnapKV.
@@ A.1.6 Quest
We implement Quest (Tang et al., 2024), which applies dynamic sparse attention during the decoding phase at the page level. Based on our ablations (Figure 6), we used a page size of 16 tokens. We represent pages by their minimum and maximum key values to enable efficient similarity computation with queries. At each decoding step, we select the most relevant pages based on query-page similarity scores, always including the page containing the current token. We control sparsity by setting the ‘token_budget‘ (number of tokens selected per step) to achieve the target compression ratio.
Table 6: Pattern parameters for different sequence lengths and compression ratios
Pattern Parameter Sequence Length Values for Different Compression Ratios
@@ A.2 Task Details
This section provides further details on the nine evaluation tasks used in our experiments. Tasks are grouped into Question Answering, synthetic tasks from RULER (Hsieh et al., 2024), and our Story tasks. We specify key hyperparameters, evaluation metrics, and characterise each task along the axes of Scope (Low vs. High) and Dispersion (Low vs. High), as defined in Table 2. Scope refers to the amount of information required, while Dispersion indicates how difficult it is to locate the relevant information within the context.
@@ A.2.1 Question Answering (QA)
We use SQuAD (Rajpurkar et al., 2018) from RULER and two other QA datasets selected for minimal data contamination (Li et al., 2024a): QuALITY (Pang et al., 2022) and ToeflQA (Tseng et al., 2016).
@@ A.2.2 Synthetic – RULER Tasks
We use three synthetic tasks from the RULER benchmark (Hsieh et al., 2024).
@@
  • Common Word Extraction (CWE): Identify the 10 most frequent words (appearing 30 times each) among distractor words (appearing 3 times each), sampled from a vocabulary of $\sim$9,000 English words [^7] https://github.com/mrmaxguns/wonderwordsmodule. Evaluated using Intersection-over-Union (IoU). Requires processing the entire context to count frequencies, demanding Low Dispersion (words are presented directly as a list, not obscured within complex structures) but High Scope (all words must be processed to determine frequencies). See Section G.3.
@@ A.2.3 Semi-Synthetic – Story Tasks
These tasks use procedurally generated multi-chapter narratives that scale with sequence length. Each chapter follows a schema involving travel, dialogue, and item transactions. See Appendix F for an example narrative.
@@
Appendix B Overview of Scaling Laws Studies
Scaling laws research aims to model the relationship between model performance and various factors such as model size, training data, and inference budget. As Li et al. (2025a) comprehensively discuss, these studies typically involve regressing performance metrics onto those factors using data from systematic experiments. We briefly outline key aspects of this field to situate our contribution within the broader scaling laws landscape.
Performance metrics
The choice of target metric depends on the controlled factors. For studies varying model size or data quantity, language modelling metrics like perplexity are common. For supervised models or capability-specific studies (e.g., RAG performance in Yue et al. (2025)), task accuracy is typically used. In our work, we measure performance using downstream accuracy on long-context benchmarks (QA, RULER, Story), which directly assess the capabilities we aim to model.
Controlled factors
Most scaling laws research focuses on training compute, which scales with model size and parameter count. Recent work has expanded to study model sparsity (MoE or activation sparsity), quantisation precision, and inference budgets. Our work contributes to this last category by controlling attention sparsity (via compression ratio), which directly impacts inference compute and memory requirements, alongside model size and sequence length.
Experimental approach
Scaling law data can be gathered either through controlled experiments (Kaplan et al., 2020; Hoffmann et al., 2022) or by analysing existing models across different families (Choshen et al., 2024; Ruan et al., 2024). The former approach is costly but precise, while the latter leverages public data but may be affected by undisclosed differences in training procedures. We conduct controlled experiments with the Qwen 2.5 family across a range of sparse attention algorithms, sparsity levels (compression ratios), sequence lengths, and tasks.
Fitting methodology
Common approaches to fitting scaling laws include:
  • Observing performance trends to determine the appropriate functional form (e.g., power law, exponential) (Luo et al., 2025). We adopt a log-linear form, assuming a power law relationship for numerical factors.
  • Fitting parameters using optimisation algorithms (typically L-BFGS) with robust loss functions like Huber loss (Abnar et al., 2025). We use L-BFGS with Huber loss.
  • Managing outliers through careful loss function parameterisation (Frantar et al., 2023) or selective data filtering (Li et al., 2025a). Our use of Huber loss helps mitigate outlier influence.
  • Validating on held-out examples, particularly those at the extremes of the parameter range (Yue et al., 2025). We validate by holding out the maximum values along each axis (model size, sequence length, compression ratio).
Our approach aligns with these established practices, as detailed in the main paper and Appendix C.
Appendix C Sparse Attention Scaling Laws: Formulation and Fitting Details
This section provides further details on our approach to formulating and fitting scaling laws for sparse attention, complementing the discussion in the main paper (Section 4.4). We aim to predict downstream task accuracy based on model size, sequence length, compression ratio, and task characteristics, focusing on the Qwen 2.5 model family and the best-performing sparse patterns (Vertical-Slash for prefilling, Quest for decoding).
C.1 Scaling law formulation
Inspired by prior work, particularly Yue et al. (2025), we develop a log-linear scaling law formulation suitable for modelling inference performance on long-context tasks:
Target
We transform task accuracy $a\in[0,1]$ to its logit $s=\sigma^{-1}(a)=\log(a)-\log(1-a)$ to create an unbounded regression target, suitable for linear modelling.
Model size
We model the effect of parameter count $N$ as $\alpha\log N$. Based on general observations that larger models perform better, $\alpha$ is expected to be positive.
Sequence length
We include a term $\beta\log L$, where $L$ is sequence length. As longer contexts often pose greater challenges, $\beta$ is expected to be negative.
Compression ratio
We incorporate compression ratio $C=\text{FLOPS}_{\text{dense}}/\text{FLOPS}_{\text{sparse}}$ as $\gamma\log C$. Since higher compression (more sparsity) typically degrades performance, $\gamma$ is expected to be negative.
Task-specific effects
We include a task-specific intercept parameter $\delta_{task}$ to account for inherent differences in task difficulty and sensitivity to the other factors.
Shared Intercept
We include a shared intercept $\epsilon$ representing the baseline performance level.
The complete scaling law formula, as presented in the main paper, is:
$s=\alpha\log N+\beta\log L+\gamma\log C+\delta_{task}+\epsilon$ (2)
C.2 Fitting procedure
We optimise the scaling law parameters ($\alpha,\beta,\gamma,\{\delta_{task}\},\epsilon$) using the L-BFGS algorithm. To improve robustness against potential outliers in the performance data, we employ a Huber loss function with $\delta=1$. As noted in the main paper, we perform three separate fitting runs for extrapolation analysis, each time holding out data points corresponding to the maximum value of one axis (model size, sequence length, or compression ratio). We evaluated a Mean Squared Error loss as well, finding similar results and consistent patterns regarding which tasks and extrapolation axes were more challenging to model accurately.
C.3 Scaling law accuracy (MAE)
As discussed in the main paper, we evaluate the extrapolation capabilities of our fitted scaling laws using Spearman’s $\rho$ (Table 4) and Mean Absolute Error (MAE) on the held-out test sets. Table 7 reports the MAE between the predicted accuracy (transformed back from logit space) and the true accuracy for each task across the different extrapolation axes (holding out max model size, sequence length, or compression ratio).
Lower MAE indicates better predictive accuracy. While the overall fit is strong (high $R^{2}$, see Table 3), the MAE values highlight specific task–axis combinations where extrapolation is more challenging. For instance, extrapolating the accuracy when holding out data for the highest compression ratio yields a relatively high MAE in Story Retrieval during prefilling (0.41) and Ruler NIAH during decoding (0.29). However, no single task or extrapolation axis proves universally difficult across both prefilling and decoding. An interesting observation is the tendency for task–axis pairs with high MAE in prefilling to also exhibit high MAE in decoding (e.g., Story Retrieval on sequence length extrapolation), suggesting that the inherent sensitivity of certain tasks to specific factors persists regardless of whether sparsity is applied during prefilling or decoding.
Table 7: Mean Absolute Error (MAE) between predicted and true accuracy by task when extrapolating along different axes (holding out max value). Lower is better.
QA Ruler Story
QuALITY SQuAD TOEFL CWE NIAH VT Filtering Multi-hop Retrieval
Prefill Model Size 0.16 0.11 0.06 0.11 0.02 0.22 0.24 0.08 0.07
Sequence Length 0.12 0.08 0.10 0.23 0.16 0.06 0.04 0.06 0.20
Compression Ratio 0.13 0.10 0.12 0.18 0.24 0.10 0.02 0.05 0.41
Decode Model Size 0.19 0.15 0.08 0.07 0.04 0.29 0.19 0.19 0.02
Sequence Length 0.16 0.12 0.17 0.14 0.19 0.08 0.04 0.07 0.24
Compression Ratio 0.10 0.10 0.08 0.12 0.29 0.10 0.07 0.05 0.15
Appendix D FLOPS breakdown
Figure 12: Computational cost of Qwen models (7B to 72B) across sequence lengths from 16K to 128K. Top: Prefilling FLOPS (TFLOPS) showing both attention (solid lines) and non-attention components (dashed lines). Bottom: Decoding FLOPS (GFLOPS) with similar decomposition. Where the lines cross shows the sequence length at which attention FLOPS equal non-attention FLOPS.
We estimate the computational cost of models from Qwen 2.5 family using a component-wise approach:
$\text{FLOPS}_{\text{total}}$ $=\text{FLOPS}_{\text{embedding}}+\text{FLOPS}_{\text{attention}}+\text{FLOPS}_{\text{mlp}}+\text{FLOPS}_{\text{logits}}$ (3)

### v2

@@ 5 Conclusions
This study provides the largest-scale empirical analysis of training-free sparse attention to date, covering three model families (Qwen 2.5, Llama 3.1, Gemma 3), model scales (4B–72B parameters), sequence lengths (16K–128K tokens), sparsity levels up to 0.95 (i.e., $1/20$ attention budget), and nine diverse long-sequence tasks. We organise the rapidly evolving landscape of sparse attention methods into a taxonomy along four design axes and introduce novel benchmarks consisting of natural texts that are fully controllable yet challenging. Our analysis yields three key insights.
Evidence of effectiveness. Sparse attention enables larger models to outperform smaller dense ones at equivalent computational cost, improving the Pareto frontier. Thus, sparsity becomes crucial for optimal LLM scaling.
Practical deployment guidance. Method selection should be task-aware: fine-grained token selection (e.g., Vertical-Slash) excels at retrieval, chunk-based methods (e.g., Block-Sparse) suit reasoning and aggregation, and Quest provides robust decoding across most scenarios.
Design recommendations. Longer sequences tolerate higher sparsity while maintaining accuracy. This suggests that fixed-budget methods deployed in production are suboptimal; future designs should adapt sparsity levels to sequence length, possibly growing the token budget sublinearly.
Limitations
First, we evaluate only training-free sparse attention methods. Training-based approaches could reduce train-inference mismatch, but require substantial computational resources and access to proprietary training data.
Second, our experimental coverage, while extensive, is bounded. We evaluate three model families (Qwen 2.5, Llama 3.1, Gemma 3) that met our methodological requirements for controlled scaling experiments with native long-context support; other families may exhibit different behaviour. We test only instruction-tuned models; reasoning models with extended chain-of-thought capabilities (e.g., o1, DeepSeek-R1) may have different attention patterns and sparsity tolerance. Our nine tasks, though selected to span diverse dispersion levels, processing scopes, and data naturalness, do not exhaustively cover all long-context scenarios—open-ended tasks like summarisation were excluded due to unreliable automated metrics. Additionally, experiments at 128k tokens are limited due to low baseline performance and lack of robustness across models; more conclusive evidence on how sequence length affects sparse attention scaling requires stronger long-context models.
Third, we report hardware-agnostic computational costs (FLOPs and memory access) rather than wall-clock timings. Actual speedups depend on hardware, batch size, and implementation quality, which vary across deployment environments.
Fourth, we do not investigate interactions between sparse attention and other model efficiency techniques such as quantisation, weight pruning, or mixture-of-experts sparsity. These methods are often combined in practice, and their joint effects on attention sparsity tolerance remain unexplored.
@@ A.1 Implementation Details
This section provides supplementary details on the sparse attention patterns evaluated, focusing on hyperparameter tuning and specific configurations used to achieve target sparsity levels. We tuned hyperparameters for each pattern using ablation studies on the Qwen-7B model with a 16K sequence length across all tasks, varying sparsity from 0 to 0.9. Our main experiments evaluated performance at sparsity levels 0.33, 0.5, 0.6, 0.7, 0.8, 0.87, 0.9, 0.93, 0.95 (corresponding to attention budgets $1/1.5$, $1/2$, $1/2.5$, $1/3.33$, $1/5$, $1/7.5$, $1/10$, $1/15$, $1/20$), using linear interpolation for intermediate values where necessary. Table 3 summarizes the final parameters we used for each pattern, sequence length, and sparsity level. In total we evaluated 7065 configurations with 100 samples per configuration. We used approximately 4 compute nodes with 8 H100 GPUs each for 21 days.
Figure 4: Visualization of sparse attention patterns. Block-Sparse and Vertical-Slash operate during prefilling (showing query-key attention matrix), while SnapKV and Quest operate during decoding (showing decoded tokens attending to KV cache positions). Colors indicate different sparsity levels from 0.5 (red) to 0.9 (blue). The black vertical lines in SnapKV and Quest mark the prefill/decode boundary.
@@ A.1.1 Block-Sparse Attention
We implement block-sparse attention by dividing the attention matrix into fixed-size blocks. The original implementation is available under the MIT license. Based on our ablation studies (Figure 6), we selected a block size of 16x16, as smaller blocks consistently yielded better performance. To achieve a target sparsity level, we select the top-$k$ key blocks for each query block, where $k$ is determined via binary search. We always preserve attention sinks (the first key block) and local context (diagonal key blocks corresponding to the query block).
@@ A.1.2 Vertical-Slash Pattern
We implement the Vertical-Slash pattern (Jiang et al., 2024), available under the MIT license, by allocating a uniform budget to global (vertical columns) and local (slash diagonals) attention components. We select the most important verticals and slashes by approximating attention scores using a limited window of recent query tokens. Our ablation studies (Figure 11) revealed task-dependent optimal approximation window sizes: 512 tokens for retrieval-heavy tasks (Ruler NIAH, Story Retrieval) and 256 tokens for other tasks. This observation correlates with the typical query lengths for these tasks (see Table 2). We consistently preserve the first 4 (prefix) and the most recent 64 (local) tokens. To achieve target sparsity levels, we compute the required number of verticals and slashes based on collected attention statistics for each sequence length.
@@ A.1.3 FlexPrefill
We implement FlexPrefill (Lai et al., 2025), available under the Apache-2.0 license, which enhances Vertical-Slash by introducing dynamic budget allocation per layer and head, controlled by a coverage parameter $\alpha$ and a minimum budget (min_budget). We set $\tau=0$ in our experiments, hence disabling Query-Aware attention. This choice stemmed from two key considerations: first, our preliminary tests indicated no significant performance gains from enabling it, aligning with the findings reported in the original work; second, this setting isolates the dynamic budget allocation mechanism, allowing us to specifically evaluate its impact compared to the fixed allocation used in the Vertical-Slash pattern. We employ the same task-dependent approximation windows (256 / 512 tokens) and critical token preservation strategy (first 4 prefix, most recent 64 local) as in our Vertical-Slash implementation. Our ablations (Figure 8) indicated that setting min_budget to 512 significantly improved performance, suggesting the importance of maintaining a minimum level of connectivity during prefilling. We achieved target compression ratios by selecting the appropriate $\alpha$ based on attention statistics while keeping min_budget fixed at 512. For high compression ratios where dynamic allocation proved less effective, we set $\alpha=0$, effectively reverting to a uniform allocation of Vertical-Slash.
@@ A.1.4 SnapKV
We implement SnapKV (Li et al., 2024b), available under the CC-BY 4.0 license, by compressing the Key-Value (KV) cache after the prefilling stage and applying a uniform token budget across all heads for the subsequent decoding phase. We predict token importance for decoding by computing attention scores using a window of recent query tokens (approximation window). Our ablations showed an optimal approximation window size of 256 tokens (Figure 10), with no significant task dependency observed, unlike Vertical-Slash and FlexPrefill. We smooth the calculated token importance scores using 1D average pooling with a kernel size of 21 (chosen based on Figure 10). We always preserve the first 4 and the most recent 128 tokens. We control sparsity by setting the ‘token_capacity‘ (token limit per head) to achieve the target sparsity level.
@@ A.1.5 Ada-SnapKV
We implement Ada-SnapKV (Feng et al., 2024), available under the MIT license, which extends SnapKV by incorporating dynamic token budget allocation per head. One difference in our implementation between Ada-SnapKV and SnapKV is that we use max-aggregation (instead of averaging) across query positions and heads for score calculation; this empirically proved more effective for adaptive allocation but had no effect for uniform (SnapKV) allocation. We utilize the same smoothing kernel size (21) and critical token preservation strategy (first 4 prefix, most recent 128 local) as in our SnapKV implementation. Our ablations (Figure 8) indicated that providing each head with a minimum budget of 20% of its capacity was optimal. Performance was found to be less sensitive to minimum budget (performing well within the 10-50% range) compared to FlexPrefill’s sensitivity, but degraded sharply when approaching 100% (uniform allocation), underscoring the benefits of dynamic allocation during decoding. We control sparsity by setting ‘token_capacity‘, identically to SnapKV.
@@ A.1.6 Quest
We implement Quest (Tang et al., 2024), available under the CC-BY 4.0 license, which applies dynamic sparse attention during the decoding phase at the page level. Based on our ablations (Figure 6), we used a page size of 16 tokens. We represent pages by their minimum and maximum key values to enable efficient similarity computation with queries. At each decoding step, we select the most relevant pages based on query-page similarity scores, always including the page containing the current token. We control sparsity by setting the ‘token_budget‘ (number of tokens selected per step) to achieve the target sparsity level.
Table 3: Pattern parameters for different sequence lengths and sparsity levels. At 128k tokens, we only evaluate Vertical-Slash and Quest.
Pattern Parameter Sequence Length Values for Different Sparsity Levels
@@ A.2 Task Details
This section provides further details on the nine evaluation tasks used in our experiments. These tasks, summarized in Table 4 at the end of this subsection, are grouped into Question Answering, synthetic tasks from RULER (Hsieh et al., 2024), and our Story tasks. We specify key hyperparameters, evaluation metrics, and characterize each task along the axes of Scope (Low vs. High) and Dispersion (Low vs. High). Scope refers to the amount of information required, while Dispersion indicates how difficult it is to locate the relevant information within the context.
@@ A.2.1 Question Answering (QA)
We use SQuAD (Rajpurkar et al., 2018) (CC BY-SA 4.0) from RULER (Apache-2.0) and two other QA datasets selected for minimal data contamination (Li et al., 2024a): QuALITY (Pang et al., 2022) (CC BY 4.0) and ToeflQA (Tseng et al., 2016)§§§Originally released at https://github.com/iamyuanchung/TOEFL-QA. License information is missing; a GitHub issue requesting clarification was opened on August 7, 2023, but has not received a response..
@@ A.2.2 Synthetic – RULER Tasks
We use three synthetic tasks from the RULER benchmark (Hsieh et al., 2024) (Apache-2.0).
@@
  • Common Word Extraction (CWE): Identify the 10 most frequent words (appearing 30 times each) among distractor words (appearing 3 times each), sampled from a vocabulary of $\sim$9,000 English wordshttps://github.com/mrmaxguns/wonderwordsmodule. Evaluated using Intersection-over-Union (IoU). Requires processing the entire context to count frequencies, demanding Low Dispersion (words are presented directly as a list, not obscured within complex structures) but High Scope (all words must be processed to determine frequencies). See Section G.3.
@@ A.2.3 Semi-Synthetic – Story Tasks
These tasks use procedurally generated multi-chapter narratives that scale with sequence length. Each chapter follows a schema involving travel, dialogue, and item transactions. See Appendix F for an example narrative. We release them under the CC BY 4.0 license.
@@
Task Name Description Dispersion Scope Natural
A.3 Model Details
Our choice of Qwen 2.5 as the primary model family was driven by strict methodological requirements. We needed a model family satisfying three criteria simultaneously: (1) native 128k context support, since sparse attention benefits emerge primarily at very long sequences; (2) multiple model sizes maintaining reasonable (non-random) performance across all sequence lengths with consistent training procedures on identical data to enable rigorous size-based comparisons; and (3) instruction-tuned versions for chain-of-thought evaluation to mitigate short-output evaluation bias toward sparse decoding methods with dense prefill (Yuan et al., 2024).
After evaluating available open-source families, Qwen 2.5 was the only one meeting these requirements. Other families were excluded for the following reasons:
  • Command: Different training data across sizes (8B vs 32B/104B).
  • Llama 3.1/3.2: Smaller models (1B, 3B) failed at 16k–32k sequence length on most tasks; 405B exceeded our computational budget.
  • Mistral: Multiple fine-tunes but fewer than three sizes with consistent training.
  • Phi-3: The 14B model showed unexpectedly poor long-context performance, worse than 4B according to RULER evaluations.
  • Yi: Limited to 32k sequence length.
  • GPT-OSS: Only two model sizes; additionally requires custom attention implementation with variable attention head biases, which is not supported by training-free sparse attention methods.
To broaden our scope and provide additional evidence for our findings, we also evaluate on Llama 3.1 (8B, 70B) and Gemma 3 (4B, 12B, 27B). Gemma 3 employs hybrid attention where 5 out of 6 layers use sliding window attention with a window size of 1024 tokens, while every 6th layer uses global (dense) attention. This makes Gemma 3 particularly interesting for our study: it is already heavily sparsified by design, and we apply training-free sparse attention methods only to the dense layers. This allows us to analyze whether additional sparsification benefits models that already incorporate architectural sparsity.
Family Size Layers Q Heads KV Heads Huggingface
Llama 3.1 8B 32 32 8 meta-llama/Llama-3.1-8B-Instruct
70B 80 64 8 meta-llama/Llama-3.1-70B-Instruct
Qwen 2.5 7B 28 28 4 Qwen/Qwen2.5-7B-Instruct
14B 48 40 8 Qwen/Qwen2.5-14B-Instruct
32B 64 40 8 Qwen/Qwen2.5-32B-Instruct
72B 80 64 8 Qwen/Qwen2.5-72B-Instruct
Gemma 3 4B 34 8 4 google/gemma-3-4b-it
12B 48 16 8 google/gemma-3-12b-it
27B 62 32 16 google/gemma-3-27b-it
Table 5: Overview of models used in the evaluation. Qwen 2.5 and Llama 3.1 officially support context lengths up to 128k tokens; however, we evaluate 128k only for Vertical-Slash and Quest. Gemma 3 supports up to 128k tokens but exhibited near-zero performance at this length across most configurations, so we evaluate Gemma 3 up to 64k only. We use 100 samples per configuration for Qwen and 50 for Llama and Gemma.
Appendix B Computational Cost Analysis
We analyze computational cost using implementation-agnostic metrics that correlate with wall-clock time under optimized implementations: FLOPs for prefilling (compute-bound) and memory transfers for decoding (memory-bound). This approach avoids confounds from implementation-specific inefficiencies while capturing the fundamental cost structure.
B.1 Cost Formulas
For prefilling, which is compute-bound, we compute total FLOPs as:
$\text{FLOPS}_{\text{prefill}}$ $=B\cdot(\text{FLOPS}_{\text{embedding}}+\text{FLOPS}_{\text{attention}}+\text{FLOPS}_{\text{mlp}}+\text{FLOPS}_{\text{logits}})$ (1)